Plz tell me the easiest method to solve this question...
tan (107n)=cos 96 + sin 96/ cos 96 - sin 96
all angles r mentioned in degrees,
we have 2 find out the value of n, which should be the smallest natural number...

Grade:

1 Answers

Rohan Maini
6 years ago
(cos96°+cos6°)/(cos96°-cos6°)=-(2cos51°cos45°)/(2sin51°sin45°)=-cot51°=tan(90°+51°)=tan141°=tan(180°+141°)=tan(321°)Therefore n=3